Two towers AB and CD are situated at a distance 'd' apart. AB is 20m  high and CD is 30m. An object of mass m is thrown from AB horizontally at 10m/s. At the same time, an object of mass 2m is thrown from top CD at an angle 60degree with the horizontal towards AB with the same magnitude of initial velocity as first. The objects collide in mid air & stick to each other.
i) Calulate the distance betweem the towers. ii)Find the position where the objects meet.

Grade:

1 Answers

siddharth gupta
9 years ago
Let us assume that the two objects meet after time t.no matter what their masses are both will experience same accln due to gravity.
let us write down the eqns of motion of particle projected from AB:
x=10t and y=1/2gt2=h
eqns of motion of particle projected from CD:
x=10cos600t and h=10sin600t+1/2gt2
when the two particles strike each other
hAB+10=hCD
1/2gt2+10=10sin600t+1/2gt2
on solving 10t=20/(3)1/2m =distance travelled by particle projected from AB and t=2/(3)1/2
i.e distance travelled by other particle is=10X1/2X2/(3)1/2=10/(3)1/2m
i.e d=10(3)1/2m.